OpenStudy (anonymous):

The legs of a right triangle measure 6 in. and 8 in. What is the length of the hypotenuse?

OpenStudy (anonymous):

To calculate the length of the hypotenuse we can use a^2+b^2=c^2. a and b are the legs and c is the hypotenuse. We get 6^2+8^2=c^2, 36+64=c^2, 100=c^2 c=10

OpenStudy (anonymous):

thnx dude